Page 3 of 12 Dickens Company produces a line of non-motorized boats. Dickens uses a job-order costing system and applies manufacturing overhead cost using direct-labor cost. The following data are for 2017 Estimated manufacturing overhead cost Actual manufacturing overhead cost Estimated direct-labor cost Actual direct-labor cost S 125,000 s 119,000 s 250,000 228,000 Inventory balances on Dec 31, 2017were as follows (before any adjustments Ending 2017 direct-labor cost in Account Work-in-process Finished goods Costs of goods sold Balance 50,700 245,050 549,250 ending balance s 20.520 s 59,280 s 148,200 S Dickens allocates any over- or under-applied overhead to the work-in-process, finished goods, and cost of goods sold accounts according to their ending balances. What is the ending balance of the three accounts at the end of the year (after under- or over-applied overhead costs a allocated)? Finished goods Work in process $71,220 $50,700 $51,000 $50,970 $145,520 S304,330 $245,050 $246,500 $245,830 $178,280 $697.450 $552.250 5352.500 S551,200 S376,200 A. B. D. Below is the information (incomplete) for Fitzgerald Corporation: Work in Process 4 Finished Goods bib 25 b/b DM DL он b 43 Calculate the total manufacturing cost (MC), cost of goods manufactured (CGM), and cost of goods COGS 283 245 265 245 255 CGM 265 245 235 A. 217 247 265 D. E. None of the above
Answer #1

3)

overhead rate =estimated overhead /estimated direct labor cost

        = 125000/250000

      = .50 or 50% of direct labor cost

overhead applied = actual direct labor cost * overhead rate

     = 228000*50%

      = 114000

overhead underapplied/(overapplied) =actual -applied

          = 119000-114000

            = 5000 underapplied

Ending balance (unadjusted)[A] weight allocation of underapplied overhead[B] Ending inventory (adjusted) [A+B]
work in process 50700

50700/845000= .06

.06*5000= 300 51000
finished goods inventory 245050 245050/845000= .29 .29*5000= 1450 246500
cost of goods sold 549250 549250/845000=.65 .65*5000= 3250 552500
845000

CORRECT OPTION IS "c"

4)TOTAL MANUFACTURING COST =DM+DL+OH

     =60+55+130

= 245

Cost of goods manufactured =b/b+ total manufacturing cost- e/b

= 10+245-20

   = 235

cost of goods sold = b/b + cost of goods manufactured- e/b

   =25+235-43

= 217

correct option is"c"